上級技術英語 1 単位 (選択)

Advanced Technical English  コインカー, パンカジ · マドゥカー · 特任助教 / 先端技術科学教育部

【授業目的】 The aim of the course is to enhance reading, writing, speaking and listening skills. Another goal is to establish a firm foundation for students to present their knowledge to others, in English.

【授業概要】 The primary focus will be the development of English skills: listening comprehension, reading, writing, and speaking. A key objective will be to increase and broaden knowledge of technical and scientific vocabulary. The final result: students should be able to produce English presentations, especially science-related presentations. The emphasis of the course content is intended to strengthen and expand both comprehension and expressive abilities. It is anticipated that this will include improving the basic aspects of clear writing: basic grammar, vocabulary, and recurring patterns of typical sentence structures.

【キーワード】 Technical English，テクノロジー， English presentations 【先行科目】『中級技術英語』 (1.0) 【関連科目】『実用技術英語』 (1.0) 【履修要件】 You should have received credit for ”Technical English.” 【履修上の注意】 This subject compulsory for Nichia STC students. If you are not a

Nichia STC student and want to attend this class, please contact a coordinator in your department for getting permission for your attendance.

【到達目標】

1. To develop vocabulary, sentence formations, and presentation principles of communication in order to establish a firm foundation for public speaking and technical presentations, in English

2. To develop three English skills: listening comprehension, speaking, reading

and writing. 【授業計画】

1. Technical Vocabulary

2. Foreign words and phrases, Prefixes

3. Accuracy in using words

4. Common spelling problems, Simple language

5. Avoiding unnecessary words, Jargon

6. Usage of Technical words: Different grammatical forms of the same word

7. Words: Synonyms, Spelling

8. Measuring units 9. Cause and effect

10. Comparing-contrasting

11. Descriptions: Stating a problem, describing situation and change

12. Choice of words

13. Graphs and diagrams

14. Expressing ideas and opinions

15. Analysis

【成績評価基準】 Grades will be based on class participation, a mid-term, presentations

and homework assignments. 【JABEE 合格】 [JABEE 合格] 【学習教育目標との関連】 [JABEE 関連] 【教科書】”Presenting Science, ” (Second Edition), 2008, Macmillan Language-

House. 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215978

【対象学生】 Nichia STC students and, possibly, a limited number of other students

are permitted to take this course. 【連絡先】

⇒ コインカー (共通講義棟 1F 国際連携教育開発センター, 088-656-7643, koi nkar@cicee.tokushima-u.ac.jp) ����

Advanced Technical English 1 unit (selection)

Pankaj Koinkar · Designated Assistant Professor / Graduate School of Advanced Technology and Science

Target⟩ The aim of the course is to enhance reading, writing, speaking and listening skills. Another goal is to establish a firm foundation for students to present their knowledge to others, in English.

Outline⟩ The primary focus will be the development of English skills: listening comprehension, reading, writing, and speaking. A key objective will be to increase and broaden knowledge of technical and scientific vocabulary. The final result: students should be able to produce English presentations, especially science-related presentations. The emphasis of the course content is intended to strengthen and expand both comprehension and expressive abilities. It is anticipated that this will include improving the basic aspects of clear writing: basic grammar, vocabulary, and recurring patterns of typical sentence structures.

�������⟩ Technical English, technology, English presentations Fundamental Lecture⟩ “Technical English”(1.0) Relational Lecture⟩ “Practical Technical English”(1.0) Requirement⟩ You should have received credit for ”Technical English.” Notice⟩ This subject compulsory for Nichia STC students. If you are not a Nichia

STC student and want to attend this class, please contact a coordinator in your department for getting permission for your attendance.

Goal⟩

1. To develop vocabulary, sentence formations, and presentation principles of communication in order to establish a firm foundation for public speaking and technical presentations, in English

2. To develop three English skills: listening comprehension, speaking, reading

and writing.

Schedule⟩

1. Technical Vocabulary

2. Foreign words and phrases, Prefixes

3. Accuracy in using words

4. Common spelling problems, Simple language

5. Avoiding unnecessary words, Jargon

6. Usage of Technical words: Different grammatical forms of the same word

7. Words: Synonyms, Spelling

8. Measuring units 9. Cause and effect

10. Comparing-contrasting

11. Descriptions: Stating a problem, describing situation and change

12. Choice of words

13. Graphs and diagrams

14. Expressing ideas and opinions

15. Analysis

Evaluation Criteria⟩ Grades will be based on class participation, a mid-term,

presentations and homework assignments. Jabee Criteria⟩ [JABEE 合格] Relation to Goal⟩ [JABEE 関連] Textbook⟩ ”Presenting Science, ” (Second Edition), 2008, Macmillan Language-

House. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215978

Student⟩ Nichia STC students and, possibly, a limited number of other students

are permitted to take this course.

Contact⟩

⇒ Koinkar (共通講義棟 1F 国際連携教育開発センター, +81-88-656-7643, koi nkar@cicee.tokushima-u.ac.jp) ����